OpenStudy (deadshot):

The length of the hypotenuse of a 45°-45°-90° triangle is 8 units. Find the exact value of the length of the leg.

OpenStudy (mathstudent55):

Use the Pythagorean theorem to write an equation. Then solve it for a. 

OpenStudy (deadshot):

8^2 - a^2 = a^2

OpenStudy (deadshot):

add a^2 to both sides, and it becomes 8^2 = a^4

OpenStudy (anonymous):

\[a = \frac{1}{\sqrt2} \times 8 =\frac{1}{\sqrt2} \times 4 \times 2\] \[=\frac{1}{\sqrt2} \times 4 \times \sqrt 2 \times \sqrt 2 = 4 \sqrt2 units\]

OpenStudy (deadshot):

Thanks!

OpenStudy (mathstudent55):

\(a^2 + a^2 = 8^2\) \(2a^2 = 64\) \(a^2 = 32\) \(a = \sqrt{32} \) \(a = \sqrt{16 \cdot 2} \) \(a = \sqrt{16} \cdot \sqrt{2} \) \(a = 4\sqrt{2} \) Answer: the leg is \(4\sqrt{2}\) units long.

OpenStudy (deadshot):

Thanks!

OpenStudy (mathstudent55):

wlcm